1.Restricted hours. Alana makes $5.80 per hour working in the library. To keep her job, she must make at least $116 per week: but to keep her scholarship, she must not earn more than $145 per week. What is the range of the number of hours per wek that she may work?

2. Living comfortably. Gary has figured that he needs to take home $30,000 a year to live comfortably. If the goverment gets 24% of garys income, then what must his income be for him to live comfortably?

Answer by ptaylor(2198)


1.
Let x=number of hours per week that she may work
Then $116<=$5.80(x)<=$145 divide each term of the inequality by $5.80
20<=x<=25------------------answer
CK
$5.80*20=$116
$5.80*25=$145
2.
Let x=the income that he must have to live comfortably
Now we are told that the government will get 24% of his income or (0.24x)
If the government gets 0.24x, then Gary will have x-0.24x left for himself and this has to equal $30,000. So our equation to solve is:
x-0.24x=$30,000
0.76x=$30,000 divide both sides by 0.76
x=$39,473.68-----------------------------------ans

Hope this helps----ptaylor